Madison Corporation sells three products (M, N, and O) in the following mix: 3:1:2.

Unit price and cost data are: M N O Unit

sales price $ 12​ $ 10​ $ 11​

Unit variable costs 7​ 6​ 9​

Total fixed costs are $529,000.

The selling price per composite unit for the current sales mix (rounded to the nearest cent) is:

$35.00.

$33.00.

$46.00.

$68.00.

$ 11.00.
Answer #1

Selling price per composite unit = ($12 X 3) + ($10 X 1) + ($11 X 2) = $68

4th option.
Answer #2

To find the selling price per composite unit for the current sales mix, we need to calculate the weighted average unit price using the given sales mix percentages and unit prices.

Let:

  • P(M) = Unit price of product M = $12.00

  • P(N) = Unit price of product N = $10.00

  • P(O) = Unit price of product O = $11.00

  • W(M) = Sales mix percentage of product M = 3

  • W(N) = Sales mix percentage of product N = 1

  • W(O) = Sales mix percentage of product O = 2

The formula for calculating the weighted average unit price (selling price per composite unit) is:

Weighted Average Unit Price = (P(M) * W(M) + P(N) * W(N) + P(O) * W(O)) / (W(M) + W(N) + W(O))

Substitute the values:

Weighted Average Unit Price = ($12.00 * 3 + $10.00 * 1 + $11.00 * 2) / (3 + 1 + 2) Weighted Average Unit Price = ($36.00 + $10.00 + $22.00) / 6 Weighted Average Unit Price = $68.00 / 6 Weighted Average Unit Price ≈ $11.33

Therefore, the selling price per composite unit for the current sales mix is approximately $11.33 (rounded to the nearest cent). However, among the given options, the closest value is $11.00. So the correct answer would be:

The selling price per composite unit for the current sales mix is $11.00.

answered by: Hydra Master